OpenStudy (anonymous):

Find the derivative with respect to x of the integral from 2 to x squared of the cosine of the quantity t squared plus 1, dt

OpenStudy (anonymous):

https://gyazo.com/34ac1d12d431fd3ca6ce517b4733a203

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (tkhunny):

Why have you posted a problem with no work shown? How far have you gotten? Give us SOMETHING to go on?

jimthompson5910 (jim_thompson5910):

you'll need to use the fundamental theorem of calculus http://aleph0.clarku.edu/~ma120/FTC.jpg the second part of it. Don't forget to use the chain rule this page may help when it comes to the chain rule: http://www.sosmath.com/calculus/integ/integ03/integ03.html Look at the part that has \[\LARGE F(x) = \int_{a}^{u(x)}f(t)dt\] notice just under that is \[\LARGE F \ '(x) = f{\Huge(}u(x){\Huge )}u \ '(x)\]

OpenStudy (anonymous):

@tkhunny I wasn't sure where to start with this one @jim_thompson5910 I'll give it a go

OpenStudy (tkhunny):

Insufficient response. You have a problem section or assignment because you have the required information. You CAN'T have NO idea. That's just not right. Start SOMEWHERE - ANYWHERE. You won't break it. It's just a problem statement. We like to help. Help us do that. :-)

OpenStudy (anonymous):

@tkhunny do my thoughts count? ;p

OpenStudy (anonymous):

@jim_thompson5910 would I integrate the function, find the the value then the domain in which the value exists?

jimthompson5910 (jim_thompson5910):

that sounds like too much work

jimthompson5910 (jim_thompson5910):

the fundamental theorem ties together derivatives and integrals in a way, the two cancel each other out (sorta)

jimthompson5910 (jim_thompson5910):

did what I posted above make sense at all?

jimthompson5910 (jim_thompson5910):

the two links posted

OpenStudy (anonymous):

I got part of it. Currently trying to see how I would manipulate it for \[\int\limits_{2}^{x^2}\cos(t^2+1)\] Since they have it simpler here https://gyazo.com/9f4530e3a1e2dc1c478a40f8f353166e

jimthompson5910 (jim_thompson5910):

let's start with something a bit simpler let's say we had \[\Large \int \cos(x)dx\] what is the value of that integral?

OpenStudy (anonymous):

sin(x)

OpenStudy (anonymous):

+C

jimthompson5910 (jim_thompson5910):

yep, now let's throw in endpoints what is the value of this \[\Large \int_{0}^{\pi}\cos(x)dx\] don't simplify or evaluate with calculator, just do a substitution/replacement

OpenStudy (anonymous):

\[\sin (\pi) \]

jimthompson5910 (jim_thompson5910):

I am looking for sin(pi) - sin(0)

OpenStudy (anonymous):

- sin(0)

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

The equation writer glitched :/

jimthompson5910 (jim_thompson5910):

so imagine we are now integrating with respect to t and we now have this \[\Large \int_{0}^{x}\cos(t)dt\] what is the value of this?

OpenStudy (anonymous):

Sin(x)-sin(0)?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

now if we apply the derivative to sin(x)-sin(0), we will get back to cos(x), agreed?

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

agreed? or no?

OpenStudy (anonymous):

Yes since we would just undo it

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

let's replace the x with x^2 what is the value of this integral? \[\Large \int_{0}^{x^2}\cos(t)dt\]

OpenStudy (anonymous):

sin(x^2)-sin(0) I see what you're getting at. so in the problem it would be sin(x^4+1)-sin(5)

jimthompson5910 (jim_thompson5910):

yes, sin(x^2) - sin(0) now if we apply the derivative to `sin(x^2) - sin(0)` what do we get?

OpenStudy (anonymous):

cos(x^2) - cos(0)

jimthompson5910 (jim_thompson5910):

close, but 2 things a) the derivative of a constant is 0 b) you forgot the chain rule

jimthompson5910 (jim_thompson5910):

sin(0) is a constant

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

\[\sin(x)=u'\cos(u)\]

OpenStudy (anonymous):

So \[2x \cos(x^2)\]

jimthompson5910 (jim_thompson5910):

if you really want to be technical, you apply the chain rule every time even if it's simply x (and not x^2) y = sin(x) dy/dx = cos(x)*d/dx[x] dy/dx = cos(x)*1 dy/dx = cos(x)

jimthompson5910 (jim_thompson5910):

yes, if y = sin(x^2) - sin(0), then dy/dx = 2x*cos(x^2)

jimthompson5910 (jim_thompson5910):

so that's effectively how that second link I posted is able to say \[\LARGE F(x) = \int_{a}^{u(x)}f(t)dt\] \[\LARGE F \ '(x) = f{\Huge (}u(x){\Huge )}u \ '(x)\]

OpenStudy (tkhunny):

Thoughts? Absolutely! Talk to us! That will work. :-)

OpenStudy (anonymous):

Ok so I have

OpenStudy (anonymous):

\[2x^2 \sin(x^4+1) - 2x^2\sin(5)\] @jim_thompson5910 would I just take the derivative?

OpenStudy (tkhunny):

You don't have to tag in every post. Anyone contributing to the thread is likely to get a notification anytime you post.

OpenStudy (anonymous):

Tag is to show who it was directed at

OpenStudy (anonymous):

@jim_thompson5910 I got D for the final answer. Mind checking?

jimthompson5910 (jim_thompson5910):

In this case, the u(x) function is x^2 what is u ' (x) equal to?

jimthompson5910 (jim_thompson5910):

I think you took the derivative of x^4 when you shouldn't have done so

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

I took the derivative of t^2 which is 2t then plugged in x^2

OpenStudy (anonymous):

Unless you're talking about the second part.

jimthompson5910 (jim_thompson5910):

So, \[\LARGE F(x) = \int_{a}^{u(x)}f(t)dt\] \[\LARGE F(x) = \int_{2}^{x^2}\cos(t^2+1)dt\] we have a = 2, u(x) = x^2 and f(t) = cos(t^2+1) \[\LARGE F \ '(x) = f{\Huge (}u(x){\Huge )}u \ '(x)\] \[\LARGE F \ '(x) = f{\Huge (}x^2{\Huge )}*2x\] \[\LARGE F \ '(x) = 2x*f{\Huge (}x^2{\Huge )}\] \[\LARGE F \ '(x) = 2x*\cos\left((x^2)^2+1\right)\] \[\LARGE F \ '(x) = 2x*\cos\left(x^4+1\right)\]

OpenStudy (anonymous):

Where I did take the derivative of x^4

jimthompson5910 (jim_thompson5910):

when you said it was D

OpenStudy (anonymous):

Yeah, so i should've taken it of x^2 instead. I see my mistake. Thanks for pointing it out.

jimthompson5910 (jim_thompson5910):

no problem